### v1

# On the Exteneded Painlevé Property of Riccati Hiearchies

Abstract
This article presents a new expression for Riccati chains using Bell polynomials. The quasi-homogeneous properties of these chains are explored, and the Kovalevskaya exponents are analyzed. Additionally, we demonstrate that Riccati chains exhibit the extended Painlevé property, employing the method of weighted projective space.
1 Introduction
The Riccati equation is a fundamental nonlinear differential equation that serves as the simplest nonlinear analog of the second-order linear equation. Its higher-order generalizations, known as Riccati chains, arise naturally when taking successive covariant derivatives with respect to the logarithmic derivative $v=\frac{f_{x}}{f}$ of a projective field $f$. These chains form a hierarchy that links linear, nonlinear, and integrable systems through the operator $L=\frac{\partial}{\partial x}+v$ [1, 2].
Riccati chains exhibit properties crucial for integrability, such as quasi-homogeneity. A vector field $f=(f_{1},\dots,f_{m})$ is quasi-homogeneous if it satisfies a scaling property under the action of a C*-group, a feature that influences both the dynamics and integrability of the system. This quasi-homogeneous structure is important in the study of the Painlevé property, which ensures that the solutions of the equations do not have movable singularities [3, 8]. In this work, we investigate the extended Painlevé property for Riccati chains using the method of weighted projective space, which is a powerful tool for studying the integrability of differential systems [5, 6, 7].
To facilitate the study of these chains, we express the Riccati chain equations using Bell polynomials, which provide a compact and efficient representation. Bell polynomials capture the recursive structure of Riccati chains, making the analysis of their singularities more tractable. Additionally, we compute the Kovalevskaya exponents, which are essential for classifying the singularities of these systems and determining their integrability.
In this article, we demonstrate that Riccati chains satisfy the extended Painlevé property, showing that they are integrable in the extended sense. This result builds upon earlier work on the Painlevé equations and the classification of integrable systems [12, 13].
The article is structured as follows: Section 2 introduces the theoretical framework and formulation of the Riccati chain equations. Section 3 discusses the quasi-homogeneity and weight classification of Riccati chains. Section 4 presents the computation of Kovalevskaya exponents and the analysis of indicial loci. Section 5 proves that Riccati chains possess the extended Painlevé property using the weighted projective space method. We conclude in Section 6 with a discussion of the implications of our results and potential future directions.
2 Theoretical Setting
In this section, we present the theoretical framework for the Riccati chain, focusing on its recursive structure and the use of Bell polynomials for a compact representation.
Definition 2.1.
Let $L$ be the first-order differential operator
$L=\frac{d}{dz}+cu(z),\quad c\in\mathbb{R}^{*},$
where $u(z)$ is a potential function. The $m$-order Riccati chain equation is a generalization of the standard Riccati equation and can be written as:
$L^{m}u(z)+\sum_{j=1}^{n}\alpha_{j}(z)L^{j-1}u(z)+\alpha_{0}(z)=0,$
where $m\in\mathbb{N}$, and the coefficient functions $\alpha_{j}(z)$ may depend on $u(z)$ and its derivatives.
Remark 2.2.
The operator $L^{m}u$ expands recursively as
$L^{k}u=\left(\frac{d}{dz}+u\right)L^{k-1}u,$
generating a polynomial in $u$ and its derivatives.
The recursive nature of this system suggests a way to express the Riccati chain efficiently using Bell polynomials, which offer a systematic way to represent the terms in the Riccati chain. The following definitions and propositions provide the necessary tools for this representation.
Definition 2.3 (Exponential partial Bell polynomial).
For integers $m\geq k\geq 1$, the exponential partial Bell polynomial $B_{m,k}(x_{1},x_{2},\dots,x_{m-k+1})$ is defined by the generating function
$\frac{1}{k!}\!\left(\sum_{n=1}^{\infty}x_{n}\,\frac{t^{n}}{n!}\right)^{\!k}=\sum_{m=k}^{\infty}B_{m,k}(x_{1},x_{2},\dots,x_{m-k+1})\,\frac{t^{m}}{m!}.$
Equivalently, let
$\mathcal{J}_{m,k}=\Bigl\{(j_{1},\dots,j_{m-k+1})\in\mathbb{N}_{0}^{\,m-k+1}\;\Big|\;\begin{array}[]{l}j_{1}+\cdots+j_{m-k+1}=k,\\[3.0pt] 1j_{1}+2j_{2}+\cdots+(m-k+1)j_{m-k+1}=m\end{array}\Bigr\}.$
Then
$B_{m,k}(x_{1},\dots,x_{m-k+1})=\sum_{(j_{1},\dots,j_{m-k+1})\in\mathcal{J}_{m,k}}\frac{m!}{j_{1}!\,j_{2}!\cdots j_{m-k+1}!}\left(\frac{x_{1}}{1!}\right)^{\!j_{1}}\left(\frac{x_{2}}{2!}\right)^{\!j_{2}}\!\cdots\left(\frac{x_{m-k+1}}{(m-k+1)!}\right)^{\!j_{m-k+1}}.$
Each $(j_{1},\dots,j_{m-k+1})\in\mathcal{J}_{m,k}$ corresponds to a way of partitioning $m$ into $k$ parts, where $j_{i}$ denotes the number of parts of size $i$.
Example 2.4.
Some examples of the Bell polynomials are:
$B_{1,1}$ $=x_{1},$
$B_{2,1}$ $=x_{2},\quad B_{2,2}=x_{1}^{2},$
$B_{3,1}$ $=x_{3},\quad B_{3,2}=3x_{1}x_{2},\quad B_{3,3}=x_{1}^{3},$
$B_{4,1}$ $=x_{4},\quad B_{4,2}=4x_{1}x_{3}+3x_{2}^{2},\quad B_{4,3}=6x_{1}^{2}x_{2},\quad B_{4,4}=x_{1}^{4}.$
These polynomials provide the building blocks for the Bell-polynomial expansion of the Riccati chain.
Proposition 2.5 (Bell–polynomial form of the Riccati chain).
Fix $c\in\mathbb{R}$ and write $L:=\partial_{z}+c\,u$. For $m\in\mathbb{N}$ and smooth coefficients $\alpha_{j}(z)$, the $m$-th Riccati–chain equation
$L^{\,m}u(z)\;+\;\sum_{j=1}^{m}\alpha_{j}(z)\,L^{\,j-1}u(z)\;+\;\alpha_{0}(z)\;=\;0$
is equivalent to
$\sum_{k=1}^{m+1}c^{\,k-1}\,B_{m+1,k}\!\big(u,u^{\prime},\dots,u^{(m)}\big)\;+\;\sum_{j=1}^{m}\alpha_{j}(z)\!\sum_{k=1}^{j}c^{\,k-1}\,B_{j,k}\!\big(u,u^{\prime},\dots,u^{(j-1)}\big)\;+\;\alpha_{0}(z)\;=\;0,$
where $B_{n,k}$ are the exponential partial Bell polynomials.
Proof.
The standard identity for iterates of $L=\partial_{z}+c\,u$ gives
$L^{\,m}u=\sum_{k=1}^{m+1}c^{\,k-1}\,B_{m+1,k}\!\big(u,u^{\prime},\dots,u^{(m)}\big),$
which follows by induction from the recursion $L^{\,m+1}u=(\partial_{z}+cu)\,L^{\,m}u$ and the Bell–polynomial recurrence
$B_{n+1,k}(_{1},\dots,x_{n+1})=\sum_{r=0}^{n-k+1}\binom{n}{r}\,x_{r+1}\,B_{n-r,k-1}(x_{1},\dots,x_{n-r}).$
Applying the same expansion to each $L^{\,j-1}u$ yields
$L^{\,j-1}u=\sum_{k=1}^{j}c^{\,k-1}B_{j,k}\!\big(u,u^{\prime},\dots,u^{(j-1)}\big).$
Substituting these two identities into the Riccati–chain equation gives the stated Bell–polynomial form. ∎
Example 2.6 (Case $m=3$).
With $L=\partial_{z}+c\,u$, the Bell–polynomial expansion gives
$L^{3}u=\sum_{k=1}^{4}c^{k-1}B_{4,k}(u,u^{\prime},u^{\prime\prime},u^{\prime\prime\prime})=u^{\prime\prime\prime}+4c\,u\,u^{\prime\prime}+3c\,(u^{\prime})^{2}+6c^{2}u^{2}u^{\prime}+c^{3}u^{4},$
since
$B_{4,1}=u^{\prime\prime\prime},\quad B_{4,2}=4u\,u^{\prime\prime}+3(u^{\prime})^{2},\quad B_{4,3}=6u^{2}u^{\prime},\quad B_{4,4}=u^{4}.$
Similarly,
$L^{2}u=\sum_{k=1}^{3}c^{k-1}B_{3,k}(u,u^{\prime},u^{\prime\prime})=u^{\prime\prime}+3c\,u\,u^{\prime}+c^{2}u^{3},\qquad Lu=u^{\prime}+c\,u^{2}.$
Substituting into
$L^{3}u+\alpha_{3}(z)\,L^{2}u+\alpha_{2}(z)\,Lu+\alpha_{1}(z)\,u+\alpha_{0}(z)=0$
yields
$u^{\prime\prime\prime}$ $+\big(\alpha_{3}(z)+4c\,u\big)u^{\prime\prime}+3c\,(u^{\prime})^{2}+\big(6c^{2}u^{2}+3c\,\alpha_{3}(z)\,u+\alpha_{2}(z)\big)u^{\prime}$
$\qquad+c^{3}u^{4}+c^{2}\alpha_{3}(z)u^{3}+c\,\alpha_{2}(z)u^{2}+\alpha_{1}(z)u+\alpha_{0}(z)\;=0,$
which matches the $m=3$ member of the Riccati chain.
3 Quasi-homogeneity and Weights
Let us consider the system of differential equations
$\frac{dx_{i}}{dz}=f_{i}(x_{1},\cdots,x_{m},z),\quad i=1,\cdots,m,$ (3.1)
where the functions $f_{i}$ are polynomials in $(x_{1},\cdots,x_{m},z)\in\mathbb{C}^{m+1}$.
The key feature of these equations that we explore is quasi-homogeneity, a property that influences the structure and behavior of the solutions. A vector field $(f_{1},\cdots,f_{m})$ is said to be quasi-homogeneous if it satisfies the following scaling property:
$f_{i}(\lambda^{p_{1}}x_{1},\dots,\lambda^{p_{m}}x_{m},\lambda^{r}z)=\lambda^{p_{i}+1}f_{i}(x_{1},\dots,x_{m},z),$
for any $\lambda\in\mathbb{C}$ and $i=1,\dots,m$, where $(p_{1},\cdots,p_{m},r)\in\mathbb{Z}^{m+1}_{>0}$ is a set of positive weights. The integers $p_{i}$ and $r$ are referred to as the weighted degrees of $x_{i}$ and $z$, respectively. This scaling property is crucial because it governs the dynamics of the system and helps classify its solutions.
The Newton polyhedron of the system (3.1) is defined as the convex hull of the union of the positive orthants in $\mathbb{R}^{m+1}$, with vertices at the exponents of the monomials in the system. The Newton diagram, which is the union of the compact faces of the Newton polyhedron, provides valuable information about the system’s solutions and their singularity structure.
If the Newton diagram consists of a single compact face, we can introduce a tuple of integers $(p_{1},\dots,p_{m},r,s)$ such that the compact face lies on the plane
$p_{1}x_{1}+\cdots+p_{m}x_{m}+rz=s$
in $\mathbb{R}^{m+1}$. In this case, the function $f_{i}$ satisfies the scaling relation:
$f_{i}(\lambda^{p_{1}}x_{1},\dots,\lambda^{p_{m}}x_{m},\lambda^{r}z)=\lambda^{s-r+p_{i}}f_{i}(x_{1},\dots,x_{m},z),$
for any $\lambda\in\mathbb{C}$.
Next, we introduce the $m$-th Riccati chain equation. By the map $x_{k}\rightarrow L^{k-1}u$, the Riccati chain can be written as a system of first-order differential equations:
$x_{k}^{\prime}$ $=x_{k+1}-cx_{1}x_{k},\quad$ $1\leq k\leq m-1,$ (3.2)
$x_{m}^{\prime}$ $=-\sum_{j=0}^{m}\alpha_{j}(z)x_{j}-cx_{1}x_{m},\quad$ $(x_{0}\equiv 1),$
which reduces to the standard Riccati equation when $m=1$:
$x_{1}^{\prime}=-\alpha_{1}(z)x_{1}-\alpha_{0}(z)+cx_{1}^{2}.$
The structure of the Riccati chain can be better understood by considering the weights of its terms. In particular, the weights of the terms $\alpha_{j}(z)$ are related to the scaling properties of the system. For example, the terms $\alpha_{j}(z)$ associated with the Riccati chain can be written as powers of $z$ with the following weights:
[…]
These weights play an essential role in understanding the quasi-homogeneous structure of the Riccati chain.
Weight Classification for the Riccati Chain: The weights of the terms $\alpha_{j}(z)$ are crucial for classifying the Riccati chain’s solutions. Specifically, for the $m$-th Riccati chain given by (3.2), the weights are given by:
$p_{k}=k\quad(1\leq k\leq m),\quad r\in\{1,\cdots,m+1\},\quad s=r+1.$
Assuming the full vector field $F=(f_{1},\dots,f_{m})$ is quasi-homogeneous of degree +1+1, the coefficient functions $\alpha_{j}(z)$ satisfy the following weight classification:
$\text{wt}(\alpha_{j})=m+1-j,\quad 0\leq j\leq m.$
In particular, if $\alpha_{j}(z)$ are polynomials in $z$, they take the form
$\alpha_{j}(z)=\begin{cases}A_{j}z^{\frac{m-j+1}{r}}&\text{if}\quad\frac{m-j+1}{r}\in\mathbb{Z},\\ 0&\text{otherwise}.\end{cases}$
This weight classification helps in understanding the structure of the Riccati chain and its integrability.
Examples for $n=1$
For $n=1$, the Riccati chain reduces to the standard Riccati equation:
$x_{1}^{\prime}=-\alpha_{1}(z)x_{1}-\alpha_{0}(z)+cx_{1}^{2}.$
The weight classification of this equation gives two types of quasi-homogeneous solutions:
$x_{1}^{\prime}=-z+cx_{1}^{2}\quad\text{(Airy type)},$
and
$x_{1}^{\prime}=-zx_{1}-z^{2}+cx_{1}^{2}\quad\text{(Hermite type)}.$
The Airy-type solution is discussed in [4], where a compactified Riccati equation of Airy type on a weighted projective space is studied.
Higher-order Examples and Chazy–Bureau Equations
For higher-order cases, the Riccati chain leads to more complex equations. For $n=3$, the Riccati chain in scalar form reads:
$u^{\prime\prime\prime}$ $+(\alpha_{3}(z)+4cu)u^{\prime\prime}+3c(u^{\prime})^{2}$
$+\left(6c^{2}u^{2}+3c\alpha_{3}(z)u+\alpha_{2}(z)\right)u^{\prime}$
$+c^{3}u^{4}+c^{2}\alpha_{3}(z)u^{3}+c\alpha_{2}(z)u^{2}+\alpha_{1}(z)u+\alpha_{0}(z)=0.$
For $\alpha_{j}(z)=0$ and $c=1$, this equation simplifies to the Chazy XII equation after an affine change of variables:
$u^{\prime\prime\prime}+4uu^{\prime\prime}+3(u^{\prime})^{2}+6u^{2}u^{\prime}+u^{4}=0.$
Similarly, the derivative of the pure second-order Riccati equation gives a special case of the Chazy IV equation:
$u^{\prime\prime\prime}+3uu^{\prime\prime}+3(u^{\prime})^{2}+3u^{2}u^{\prime}=0.$
For $n=4$, the Riccati chain yields the Fuchs-type equation:
$u^{(4)}+5uu^{\prime\prime\prime}+10u^{\prime}u^{\prime\prime}+10u^{2}u^{\prime\prime}+15u(u^{\prime})^{2}+10u^{3}u^{\prime}+u^{5}=0,$
which coincides with the Fuchs-type equation F-XVI in the Bureau classification.
These higher-order relations for the Riccati chain and their connection to Painlevé-type equations are thoroughly explored in [12], where the stabilizer set of the Virasoro orbit and the relations between these equations are studied.
4 Indicial Loci and Kovalevskaya Exponents
We begin by recalling the autonomous part of the $m$-th Riccati chain(eq 3.2), written as an $m$-dimensional first-order system:
$\frac{dx_{i}}{dz}=f_{i}^{A}(x),\quad i=1,\dots,m,$
where the functions $f_{i}^{A}(x)$ are given by
$f_{k}^{A}(x)$ $=x_{k+1}-cx_{1}x_{k},\quad$ $1\leq k\leq m-1,$ (4.1)
$f_{m}^{A}(x)$ $=-cx_{1}x_{m},$
and $c\in\mathbb{C}^{*}$ is a constant. We assign weights to the variables $x_{k}$ as follows:
$\mathrm{wt}(x_{k})=p_{k}:=k,\quad 1\leq k\leq m,$
so that the vector field $f^{A}=(f_{1}^{A},\dots,f_{m}^{A})$ is quasi-homogeneous of degree +1+1, meaning that it satisfies the scaling relation:
$f_{i}^{A}(\lambda^{p_{1}}x_{1},\dots,\lambda^{p_{m}}x_{m})=\lambda^{p_{i}+1}f_{i}^{A}(x_{1},\dots,x_{m}),\quad i=1,\dots,m,\quad\lambda\in\mathbb{C}^{*}.$
Next, we examine the formal Laurent series solutions of the truncated system of equations for small $z-z_{0}$. The solutions are of the form
$x_{i}(z)=\xi_{i}(z-z_{0})^{-p_{i}}+\cdots,$
where $p_{i}=i$ for $i=1,\dots,m$, and the leading coefficients $\xi=(\xi_{1},\dots,\xi_{m})$ satisfy the indicial equations
$p_{i}\xi_{i}=f_{i}^{A}(\xi),\quad i=1,\dots,m.$ (4.2)
Lemma 4.1.
For each integer $n\in\{1,\dots,m\}$, define
$\xi^{(n)}=\left(\xi^{(n)}_{1},\dots,\xi^{(n)}_{m}\right),\quad\xi^{(n)}_{k}=\begin{cases}\dfrac{(n)_{k}}{c},&1\leq k\leq n,\\ 0,&n<k\leq m,\end{cases}$
where $(n)_{k}:=n(n-1)\cdots(n-k+1)$ is the falling factorial. Then each $\xi^{(n)}$ satisfies the indicial equations (4.2). Moreover, these vectors $\xi^{(1)},\dots,\xi^{(m)}$ are the only nonzero solutions to the indicial equations.
Proof.
Substituting $x_{i}(z)=\xi_{i}(z-z_{0})^{-p_{i}}$ into (4.1) and comparing the leading terms, we obtain the following recursive relations:
$k\xi_{k}=\xi_{k+1}-c\xi_{1}\xi_{k},\quad 1\leq k\leq m-1,\quad-m\xi_{m}=-c\xi_{1}\xi_{m}.$
From the second equation, we have
$(c\xi_{1}-m)\xi_{m}=0.$
If $\xi_{m}\neq 0$, then $\xi_{1}=\frac{m}{c}$. Using the recursive structure of the system, we find that $\xi_{k}=\frac{(m)_{k}}{c}$ for $1\leq k\leq m$.
On the other hand, if $\xi_{m}=0$, the condition $(c\xi_{1}-(m-1))\xi_{m-1}=0$ must hold. Thus, either $\xi_{1}=\frac{m-1}{c}$ or $\xi_{m-1}=0$. If $\xi_{1}=\frac{m-1}{c}$, we obtain $\xi_{k}=\frac{(m-1)_{k}}{c}$ for $1\leq k\leq m-1$. Continuing this process, we obtain $\xi_{k}=\frac{(n)_{k}}{c}$ for some integer $n\in\{1,\dots,m\}$.
Thus, the only nonzero solutions to the indicial equations are the vectors $\xi^{(1)},\dots,\xi^{(m)}$, which correspond to the falling factorials $(n)_{k}/c$ for each $n$. ∎
Kovalevskaya matrix and principal exponents
For a given indicial locus $\xi\in{\mathbb{C}}^{m}\setminus\{0\}$, the Kovalevskaya matrix is defined as
$K(\xi):=\bigl(K_{ij}(\xi)\bigr)_{1\leq i,j\leq m},\qquad K_{ij}(\xi)=\frac{\partial f_{i}^{A}}{\partial x_{j}}(\xi)+p_{i}\delta_{ij},$ (4.3)
where $p_{i}=i$ and $\delta_{ij}$ is the Kronecker delta. Its eigenvalues are called the Kovalevskaya exponents associated with $\xi$.
Theorem 4.2.
The eigenvalues of $K(\xi^{(n)})$ are given by:
$\lambda_{k}=-n+k,\quad\text{for}\quad k=0,1,2,\dots,n-1,$
and
$\lambda_{k}=m-n+k,\quad\text{for}\quad k=n,n+1,\dots,m-1.$
Proof.
Let $K(\xi^{(n)})$ be the Kovalevskaya matrix for the indicial locus $\xi^{(n)}$. The matrix $K(\xi^{(n)})-\lambda I$ takes the form
$K(\xi^{(n)})-\lambda I=\begin{pmatrix}1-2n-\lambda&1&0&\cdots&0\\ -(n)_{2}&2-n-\lambda&1&\cdots&0\\ -(n)_{3}&0&3-n-\lambda&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&1\\ -(n)_{m}&0&\cdots&0&m-n-\lambda\end{pmatrix}$,
where $(n)_{k}=n(n-1)\cdots(n-k+1)$ is the falling factorial. For $m>n$, the terms $(n)_{k}=0$ for $k>n$.
We begin by observing the block structure of $K(\xi^{(n)})-\lambda I$. Since $(n)_{k}=0$ for $k>n$, the matrix can be written as:
$K(\xi^{(n)})-\lambda I=\begin{pmatrix}A_{n}&0\\ 0&C_{m-n}\end{pmatrix}$,
where: - $A_{n}$ is an $n\times n$ matrix, - $C_{m-n}$ is a upper triangular matrix of size $(m-n)\times(m-n)$.
The determinant of this block matrix is given by:
$\det\begin{pmatrix}A_{n}&0\\ 0&C_{m-n}\end{pmatrix}=\det(A_{n})\cdot\det(C_{m-n}).$
The matrix $C_{m-n}$ is an upper triangular matrix with diagonal entries:
$C_{m-n}=\operatorname{diag}(m-n-\lambda,m-n+1-\lambda,\dots,m-1-\lambda).$
Thus, its determinant is:
$\det(C_{m-n})=\prod_{k=n+1}^{m}(k-n-\lambda).$
The matrix $A_{n}$ is lower triangular, and its entries are:
$A_{n}=\begin{pmatrix}1-2n-\lambda&1&0&\cdots&0\\ -(n)_{2}&2-n-\lambda&1&\cdots&0\\ -(n)_{3}&0&3-n-\lambda&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&1\\ -(n)_{n}&0&\cdots&0&-\lambda\end{pmatrix}$.
$\det(A_{n})$ $=(1-2n-\lambda)\cdot\det(A_{n}^{(1,1)})+(n)_{2}\cdot\det(A_{n}^{(1,2)})$
$=(1-2n-\lambda)\prod_{k=2}^{n}(k-n-\lambda)+\sum_{j=2}^{n-1}(n)_{j}\prod_{\begin{subarray}{c}k=j+1\end{subarray}}^{n}(k-n-\lambda)+(n)_{n}$
$=(1-2n-\lambda)\prod_{k=2}^{n}(k-n-\lambda)+(n)_{2}\prod_{\begin{subarray}{c}k=2\end{subarray}}^{n-1}(k-n-\lambda)$
$=\left((1-2n-\lambda)\lambda+(n)(n-1)\right)\prod_{k=2}^{n-1}(k-n-\lambda)$
$=\prod_{\begin{subarray}{c}k=0\end{subarray}}^{n-1}(k-n-\lambda)$
Combining the results, the determinant of $K(\xi^{(n)})-\lambda I$ becomes:
$\det(K(\xi^{(n)})-\lambda I)=\left(\prod_{k=0}^{n-1}(k-n-\lambda)\right)\cdot\left(\prod_{k=n+1}^{m}(k-n-\lambda)\right).$
To find the eigenvalues, we set the determinant equal to zero:
$\prod_{k=0}^{n-1}(k-n-\lambda)\cdot\prod_{k=n+1}^{m}(k-n-\lambda)=0.$
Thus, for each $n\in\{1,\dots,m\}$, the eigenvalues $\kappa=(\kappa_{1},\dots,\kappa_{m})$ of $K(\xi^{(n)})$ follow the pattern $(-n,-n+1,\dots,-1,1,\dots,m-n)$, which completes the proof. ∎
Definition 4.3 (Principal and lower indicial loci).
We call the indicial locus $\xi^{(1)}$ principal, since its K-exponents are exactly $\{-1,1,\dots,m-1\}$, i.e. one distinguished eigenvalue 1-1 and m1m-1 strictly positive integers. The loci $\xi^{(n)}$ with $m\geq 2$ are called lower indicial loci, in analogy with the terminology of the Painlevé test: they do not carry the full set of nonnegative integer Kovalevskaya exponents.
5 Extended Painlevé Property
The goal of this section is to prove that the Riccati chain satisfies the Extended Painlevé Property (EPP), as defined in [5]. For details on weighted projective spaces, readers can refer to [5].
The Riccati hierarchy can be understood in the context of weighted projective spaces, which are spaces where the coordinates are scaled according to specific weights under a $\mathbb{C}^{*}$-action. More specifically, the weighted projective space $\mathbb{C}P^{m+2}(1,\cdots,m,m+1,m+2)$ associated with the $m$-th Riccati chain is defined by the equivalence relation on $\mathbb{C}^{4}\setminus\{0\}$:
$(x_{1},\dots,x_{m},z,\varepsilon)\sim(\lambda x_{1},\dots,\lambda^{m}x_{m},\lambda^{m}z,\lambda^{m+1}\varepsilon),$
where $\lambda\in\mathbb{C}^{*}$. This defines the weighted projective space, where each coordinate is rescaled according to a specific weight.
The orbifold structure of this space is given by:
$\mathbb{C}P^{m+1}(1,\cdots,m,r,s)=\mathbb{C}^{m+1}/\mathbb{Z}_{r}\cup\cdots\cup\mathbb{C}^{m+1}/\mathbb{Z}_{s},$
where each $\mathbb{C}^{m+1}/\mathbb{Z}_{i}$ is a local chart for the projective space. The inhomogeneous coordinates on each chart are denoted by $(x_{1},\dots,x_{m},z)$.
The transformation between the inhomogeneous coordinates $(x_{1},\dots,x_{m},z)$ on $\mathbb{C}^{m+1}/\mathbb{Z}_{s}$ and the coordinates $(X_{1},\dots,X_{j-1},X_{j+1},\dots,X_{m},Z,\varepsilon)$ on $\mathbb{C}^{m+1}/\mathbb{Z}_{j}$ is given by:
$x_{i}=X_{i}\varepsilon^{-1/s}\quad(i\neq j),\quad x_{j}=\varepsilon^{-j/s},\quad z=Z\varepsilon^{-r/s},$
where $\varepsilon$ is a local parameter for the action of $\mathbb{C}^{*}$.
The Riccati Chain System in Local Coordinates
The Riccati chain system on the local chart $\mathbb{C}^{m+1}/\mathbb{Z}_{j}$ can be written as the following system of differential equations:
$\left\{\begin{array}[]{l}\displaystyle\frac{dX_{i}}{d\varepsilon}=\frac{1}{s\varepsilon}\left(iX_{i}-j\frac{f_{i}+\varepsilon G_{i}}{f_{j}+\varepsilon G_{j}}\right),\quad i=1,\dots,m,\quad i\neq j,\\ \displaystyle\frac{dZ}{d\varepsilon}=\frac{1}{s\varepsilon}\left(rZ-\frac{p_{j}\varepsilon}{f_{j}+\varepsilon G_{j}}\right),\end{array}\right$. (5.1)
where $f_{i}=f_{i}(X_{1},\dots,X_{m},Z)$ are the components of the Riccati chain in the local coordinates. This system is rational and invariant under the $\mathbb{Z}_{j}$-action, reflecting the weighted projective space structure.
For convenience, we rewrite the system (5.1) as an autonomous vector field in the form:
$\left\{\begin{array}[]{l}\displaystyle\frac{dX_{i}}{dt}=iX_{i}-j\frac{X_{i+1}-cX_{1}X_{k}}{X_{j+1}-cX_{1}},\quad(i=1,\dots,m;i\neq j),\\ \displaystyle\frac{dZ}{dt}=rZ-\frac{j\varepsilon}{X_{j+1}-cX_{1}},\\ \displaystyle\frac{d\varepsilon}{dt}=s\varepsilon,\end{array}\right$. (5.2)
where the new independent variable $t$ parametrizes the integral curves of the system. This formulation makes it easier to analyze the behavior of the Riccati chain in the context of weighted projective spaces.
The following lemma describes the fixed points of the vector field in (5.2).
Lemma 5.1.
The point
$\left(X_{1},\dots,X_{m},Z,\varepsilon\right)=\begin{cases}\left(\frac{(n)_{1}}{c}\left(\frac{(n)_{j}}{c}\right)^{-1/j},\dots,\frac{(n)_{j-1}}{c}\left(\frac{(n)_{j}}{c}\right)^{-(j-1)},0,\dots,0,0\right),&\text{for }j\leq n,\\ (0,\dots,0),&\text{for }j>n,\end{cases}$ (5.3)
is a fixed point of the vector field in (5.2).
Lemma 5.2.
The eigenvalues of the Jacobi matrix of the vector field (5.2) at the fixed point (5.3) are given by $r$, $s$, and the Kovalevskaya exponents of the system (4.1), excluding the trivial exponent 1-1. This result has been established in [5].
We now state the main result of this section:
Theorem 5.3.
The Riccati chain satisfies the Extended Painlevé Property (EPP) in the sense that it admits an mk+1m-k+1-parameter family of convergent Laurent series solutions, where the set $\{a_{i,n}\}$ includes mkm-k free parameters.
Proof.
For each indicial locus $(\xi^{(n)})$, it corresponds to n1n-1 negative eigenvalues, mnm-n positive eigenvalues, and the trivial exponent 1-1. By Lemma 4.2, the eigenvalues of the Jacobi matrix at the fixed point are given by $\kappa$, and the parameters $r$ and $s$.
By translation, the system can be decomposed into an mn+2m-n+2-dimensional unstable manifold and an n1n-1-dimensional stable manifold. We perform a suitable linear transformation, which allows us to rewrite the system in the form:
$\left\{\begin{array}[]{l}\displaystyle\frac{d\mbox{$X$}_{u}}{dt}=J_{u}\mbox{$X$}_{u}+F_{u}(\mbox{$X$}_{u},\mbox{$X$}_{s}),\quad\mbox{$X$}_{u}\in\mathbb{C}^{m-n+2},\\ \displaystyle\frac{d\mbox{$X$}_{s}}{dt}=J_{s}\mbox{$X$}_{s}+F_{s}(\mbox{$X$}_{u},\mbox{$X$}_{s}),\quad\mbox{$X$}_{s}\in\mathbb{C}^{n-1}.\end{array}\right$.
where $J_{u}$ and $J_{s}$ are the Jacobian matrices corresponding to the unstable and stable manifolds, respectively, with the real parts of the eigenvalues of $J_{u}$ being positive and those of $J_{s}$ being negative.
By the unstable manifold theorem, there exists a local analytic function $\varphi$ such that $\varphi(0)=D\varphi(0)=0$, and the set $(\mbox{$X$}_{u},\varphi(\mbox{$X$}_{u}))$ forms the unstable manifold. The dynamics on the unstable manifold can then be expressed as:
$\frac{d\mbox{$X$}_{u}}{dt}=J_{u}\mbox{$X$}_{u}+F_{u}(\mbox{$X$}_{u},\varphi(\mbox{$X$}_{u})).$
Note that $J_{u}$ is semisimple and has distinct positive integer eigenvalues. Since for $j>n$, the fixed point is located at the origin, $(X_{1},\dots,X_{m},Z,\varepsilon)=(0,\dots,0,0)$, the fixed point lies at the origin.
Furthermore, since the system does not contain resonant terms in the Poincaré domain, we can apply Poincaré’s linearization theory to linearize the unstable manifold, completing the proof that the Riccati chain satisfies the Extended Painlevé Property. ∎
Acknowledgments. The author wishes to express sincere gratitude to Professor Hayato Chiba for his valuable guidance and instruction.

### v3

@@
-On the Exteneded Painlevé Property of Riccati Hiearchies
+Kovalevskaya exponents of the Riccati hierarchy
 Abstract
-This article presents a new expression for Riccati chains using Bell polynomials.
-The quasi-homogeneous properties of these chains are explored, and the Kovalevskaya exponents are analyzed.
-Additionally, we demonstrate that Riccati chains exhibit the extended Painlevé property, employing the method of weighted projective space.
+We study the Kovalevskaya exponents of the Riccati hierarchy, deriving the general indicial loci and exponents.
+A unique commuting vector field is identified, and the general solution is obtained via symmetry reduction.
+Finally, Laurent series expansions in annular regions are analyzed to reveal the singular structure of the solutions.
 1 Introduction
-The Riccati equation is a fundamental nonlinear differential equation that serves as the simplest nonlinear analog of the second-order linear equation.
-Its higher-order generalizations, known as Riccati chains, arise naturally when taking successive covariant derivatives with respect to the logarithmic derivative $v=\frac{f_{x}}{f}$ of a projective field $f$.
-These chains form a hierarchy that links linear, nonlinear, and integrable systems through the operator $L=\frac{\partial}{\partial x}+v$ [1, 2].
-Riccati chains exhibit properties crucial for integrability, such as quasi-homogeneity.
-A vector field $f=(f_{1},\dots,f_{m})$ is quasi-homogeneous if it satisfies a scaling property under the action of a C*-group, a feature that influences both the dynamics and integrability of the system.
-This quasi-homogeneous structure is important in the study of the Painlevé property, which ensures that the solutions of the equations do not have movable singularities [3, 8].
-In this work, we investigate the extended Painlevé property for Riccati chains using the method of weighted projective space, which is a powerful tool for studying the integrability of differential systems [5, 6, 7].
-To facilitate the study of these chains, we express the Riccati chain equations using Bell polynomials, which provide a compact and efficient representation.
-Bell polynomials capture the recursive structure of Riccati chains, making the analysis of their singularities more tractable.
-Additionally, we compute the Kovalevskaya exponents, which are essential for classifying the singularities of these systems and determining their integrability.
-In this article, we demonstrate that Riccati chains satisfy the extended Painlevé property, showing that they are integrable in the extended sense.
-This result builds upon earlier work on the Painlevé equations and the classification of integrable systems [12, 13].
-The article is structured as follows: Section 2 introduces the theoretical framework and formulation of the Riccati chain equations.
-Section 3 discusses the quasi-homogeneity and weight classification of Riccati chains.
-Section 4 presents the computation of Kovalevskaya exponents and the analysis of indicial loci.
-Section 5 proves that Riccati chains possess the extended Painlevé property using the weighted projective space method.
-We conclude in Section 6 with a discussion of the implications of our results and potential future directions.
-2 Theoretical Setting
-In this section, we present the theoretical framework for the Riccati chain, focusing on its recursive structure and the use of Bell polynomials for a compact representation.
+The study of Kovalevskaya exponents is central to understanding the integrability and singularities of nonlinear differential equations [10, 12, 15], particularly in the context of quasi-homogeneous systems [3].
+These exponents offer a way to analyze the behavior of solutions to integrable systems, helping to classify and understand the singularities of their Laurent series solutions.
+If we assume the existence of a hierarchy of integrable equations, it is equivalent to say that we have a family of commuting flows.
+In [16], we studied the commuting vector field under quasi-homogeneity settings.
+Motivated by this study of commuting flows in quasi-homogeneous systems and their connection to Painlevé equations, we specifically focus on the Kovalevskaya exponents of the Riccati hierarchy in this paper.
+The Riccati equation, a fundamental nonlinear differential equation, has higher-order generalizations known as Riccati chains.
+These chains naturally arise in integrability theory when taking successive covariant derivatives of a projective field, and they provide a rich structure that links linear, nonlinear, and integrable systems.
+The Kovalevskaya exponents of the Riccati hierarchy are crucial for determining the singular structure of these systems.
+The key findings of this paper are as follows:
+  • Indicial Loci and Kovalevskaya Exponents: We derive the general indicial loci and Kovalevskaya exponents for the Riccati hierarchy.
+These exponents are determined by a specific recursive structure, with the Kovalevskaya exponents revealing the free parameters at each stage of the hierarchy.
+  • Commuting Vector Field: We identify a unique commuting vector field for the Riccati hierarchy and use symmetry reduction to derive the general solution.
+This reduction provides a simplified form of the solution, aiding in a clearer understanding of the structure of the solutions.
+In the 2-dimensional case, we also apply the blow-up method commonly used in Painlevé analysis to resolve singularities, which yields the same general solution as derived through symmetry reduction.
+  • Annular Laurent Series Expansions: By analyzing Laurent series expansions in annular regions of the complex plane, we explore the singularities of the Riccati chain.
+We focus on how negative Kovalevskaya exponents influence the expansion, offering a deeper understanding of the structure of solutions in different regions.
+This work contributes to the broader field of integrable systems by providing a detailed analysis of the Kovalevskaya exponents of the Riccati hierarchy and their connection to the singularities of the system’s solutions.
+Furthermore, we demonstrate how commuting vector fields and symmetry reduction techniques can be applied to simplify the solution structure, making it more tractable for further analysis.
+Finally, the study of Laurent series expansions in various regions of the complex plane offers new insights into the singular structure of the Riccati chain and the role of negative Kovalevskaya exponents.
+2 Quasi-homogeneity of the Riccati Chain
+In this section, we introduce the theoretical framework for the Riccati chain, emphasizing its recursive structure and the central concept of quasi-homogeneity.
+Quasi-homogeneity plays a crucial role in the integrability of differential systems, particularly in systems with the Painlevé property, by governing the scaling behavior of their solutions.
 Definition 2.1.
 Let $L$ be the first-order differential operator
-$L=\frac{d}{dz}+cu(z),\quad c\in\mathbb{R}^{*},$
+$L=\frac{d}{dz}+cu(z),\quad c\in\mathbb{C}^{*},$
 where $u(z)$ is a potential function. The $m$-order Riccati chain equation is a generalization of the standard Riccati equation and can be written as:
-$L^{m}u(z)+\sum_{j=1}^{n}\alpha_{j}(z)L^{j-1}u(z)+\alpha_{0}(z)=0,$
-where $m\in\mathbb{N}$, and the coefficient functions $\alpha_{j}(z)$ may depend on $u(z)$ and its derivatives.
-Remark 2.2.
-The operator $L^{m}u$ expands recursively as
-$L^{k}u=\left(\frac{d}{dz}+u\right)L^{k-1}u,$
-generating a polynomial in $u$ and its derivatives.
-The recursive nature of this system suggests a way to express the Riccati chain efficiently using Bell polynomials, which offer a systematic way to represent the terms in the Riccati chain.
-The following definitions and propositions provide the necessary tools for this representation.
-Definition 2.3 (Exponential partial Bell polynomial).
-For integers $m\geq k\geq 1$, the exponential partial Bell polynomial $B_{m,k}(x_{1},x_{2},\dots,x_{m-k+1})$ is defined by the generating function
-$\frac{1}{k!}\!\left(\sum_{n=1}^{\infty}x_{n}\,\frac{t^{n}}{n!}\right)^{\!k}=\sum_{m=k}^{\infty}B_{m,k}(x_{1},x_{2},\dots,x_{m-k+1})\,\frac{t^{m}}{m!}.$
-Equivalently, let
-$\mathcal{J}_{m,k}=\Bigl\{(j_{1},\dots,j_{m-k+1})\in\mathbb{N}_{0}^{\,m-k+1}\;\Big|\;\begin{array}[]{l}j_{1}+\cdots+j_{m-k+1}=k,\\[3.0pt] 1j_{1}+2j_{2}+\cdots+(m-k+1)j_{m-k+1}=m\end{array}\Bigr\}.$
-Then
-$B_{m,k}(x_{1},\dots,x_{m-k+1})=\sum_{(j_{1},\dots,j_{m-k+1})\in\mathcal{J}_{m,k}}\frac{m!}{j_{1}!\,j_{2}!\cdots j_{m-k+1}!}\left(\frac{x_{1}}{1!}\right)^{\!j_{1}}\left(\frac{x_{2}}{2!}\right)^{\!j_{2}}\!\cdots\left(\frac{x_{m-k+1}}{(m-k+1)!}\right)^{\!j_{m-k+1}}.$
-Each $(j_{1},\dots,j_{m-k+1})\in\mathcal{J}_{m,k}$ corresponds to a way of partitioning $m$ into $k$ parts, where $j_{i}$ denotes the number of parts of size $i$.
-Example 2.4.
-Some examples of the Bell polynomials are:
-$B_{1,1}$ $=x_{1},$
-$B_{2,1}$ $=x_{2},\quad B_{2,2}=x_{1}^{2},$
-$B_{3,1}$ $=x_{3},\quad B_{3,2}=3x_{1}x_{2},\quad B_{3,3}=x_{1}^{3},$
-$B_{4,1}$ $=x_{4},\quad B_{4,2}=4x_{1}x_{3}+3x_{2}^{2},\quad B_{4,3}=6x_{1}^{2}x_{2},\quad B_{4,4}=x_{1}^{4}.$
-These polynomials provide the building blocks for the Bell-polynomial expansion of the Riccati chain.
-Proposition 2.5 (Bell–polynomial form of the Riccati chain).
-Fix $c\in\mathbb{R}$ and write $L:=\partial_{z}+c\,u$. For $m\in\mathbb{N}$ and smooth coefficients $\alpha_{j}(z)$, the $m$-th Riccati–chain equation
-$L^{\,m}u(z)\;+\;\sum_{j=1}^{m}\alpha_{j}(z)\,L^{\,j-1}u(z)\;+\;\alpha_{0}(z)\;=\;0$
-is equivalent to
-$\sum_{k=1}^{m+1}c^{\,k-1}\,B_{m+1,k}\!\big(u,u^{\prime},\dots,u^{(m)}\big)\;+\;\sum_{j=1}^{m}\alpha_{j}(z)\!\sum_{k=1}^{j}c^{\,k-1}\,B_{j,k}\!\big(u,u^{\prime},\dots,u^{(j-1)}\big)\;+\;\alpha_{0}(z)\;=\;0,$
-where $B_{n,k}$ are the exponential partial Bell polynomials.
-Proof.
-The standard identity for iterates of $L=\partial_{z}+c\,u$ gives
-$L^{\,m}u=\sum_{k=1}^{m+1}c^{\,k-1}\,B_{m+1,k}\!\big(u,u^{\prime},\dots,u^{(m)}\big),$
-which follows by induction from the recursion $L^{\,m+1}u=(\partial_{z}+cu)\,L^{\,m}u$ and the Bell–polynomial recurrence
-$B_{n+1,k}(_{1},\dots,x_{n+1})=\sum_{r=0}^{n-k+1}\binom{n}{r}\,x_{r+1}\,B_{n-r,k-1}(x_{1},\dots,x_{n-r}).$
-Applying the same expansion to each $L^{\,j-1}u$ yields
-$L^{\,j-1}u=\sum_{k=1}^{j}c^{\,k-1}B_{j,k}\!\big(u,u^{\prime},\dots,u^{(j-1)}\big).$
-Substituting these two identities into the Riccati–chain equation gives the stated Bell–polynomial form. ∎
-Example 2.6 (Case $m=3$).
-With $L=\partial_{z}+c\,u$, the Bell–polynomial expansion gives
-$L^{3}u=\sum_{k=1}^{4}c^{k-1}B_{4,k}(u,u^{\prime},u^{\prime\prime},u^{\prime\prime\prime})=u^{\prime\prime\prime}+4c\,u\,u^{\prime\prime}+3c\,(u^{\prime})^{2}+6c^{2}u^{2}u^{\prime}+c^{3}u^{4},$
-since
-$B_{4,1}=u^{\prime\prime\prime},\quad B_{4,2}=4u\,u^{\prime\prime}+3(u^{\prime})^{2},\quad B_{4,3}=6u^{2}u^{\prime},\quad B_{4,4}=u^{4}.$
-Similarly,
-$L^{2}u=\sum_{k=1}^{3}c^{k-1}B_{3,k}(u,u^{\prime},u^{\prime\prime})=u^{\prime\prime}+3c\,u\,u^{\prime}+c^{2}u^{3},\qquad Lu=u^{\prime}+c\,u^{2}.$
-Substituting into
-$L^{3}u+\alpha_{3}(z)\,L^{2}u+\alpha_{2}(z)\,Lu+\alpha_{1}(z)\,u+\alpha_{0}(z)=0$
-yields
-$u^{\prime\prime\prime}$ $+\big(\alpha_{3}(z)+4c\,u\big)u^{\prime\prime}+3c\,(u^{\prime})^{2}+\big(6c^{2}u^{2}+3c\,\alpha_{3}(z)\,u+\alpha_{2}(z)\big)u^{\prime}$
-$\qquad+c^{3}u^{4}+c^{2}\alpha_{3}(z)u^{3}+c\,\alpha_{2}(z)u^{2}+\alpha_{1}(z)u+\alpha_{0}(z)\;=0,$
-which matches the $m=3$ member of the Riccati chain.
-3 Quasi-homogeneity and Weights
-Let us consider the system of differential equations
-$\frac{dx_{i}}{dz}=f_{i}(x_{1},\cdots,x_{m},z),\quad i=1,\cdots,m,$ (3.1)
-where the functions $f_{i}$ are polynomials in $(x_{1},\cdots,x_{m},z)\in\mathbb{C}^{m+1}$.
-The key feature of these equations that we explore is quasi-homogeneity, a property that influences the structure and behavior of the solutions.
-A vector field $(f_{1},\cdots,f_{m})$ is said to be quasi-homogeneous if it satisfies the following scaling property:
-$f_{i}(\lambda^{p_{1}}x_{1},\dots,\lambda^{p_{m}}x_{m},\lambda^{r}z)=\lambda^{p_{i}+1}f_{i}(x_{1},\dots,x_{m},z),$
-for any $\lambda\in\mathbb{C}$ and $i=1,\dots,m$, where $(p_{1},\cdots,p_{m},r)\in\mathbb{Z}^{m+1}_{>0}$ is a set of positive weights.
-The integers $p_{i}$ and $r$ are referred to as the weighted degrees of $x_{i}$ and $z$, respectively.
-This scaling property is crucial because it governs the dynamics of the system and helps classify its solutions.
-The Newton polyhedron of the system (3.1) is defined as the convex hull of the union of the positive orthants in $\mathbb{R}^{m+1}$, with vertices at the exponents of the monomials in the system.
-The Newton diagram, which is the union of the compact faces of the Newton polyhedron, provides valuable information about the system’s solutions and their singularity structure.
-If the Newton diagram consists of a single compact face, we can introduce a tuple of integers $(p_{1},\dots,p_{m},r,s)$ such that the compact face lies on the plane
-$p_{1}x_{1}+\cdots+p_{m}x_{m}+rz=s$
-in $\mathbb{R}^{m+1}$. In this case, the function $f_{i}$ satisfies the scaling relation:
-$f_{i}(\lambda^{p_{1}}x_{1},\dots,\lambda^{p_{m}}x_{m},\lambda^{r}z)=\lambda^{s-r+p_{i}}f_{i}(x_{1},\dots,x_{m},z),$
-for any $\lambda\in\mathbb{C}$.
-Next, we introduce the $m$-th Riccati chain equation. By the map $x_{k}\rightarrow L^{k-1}u$, the Riccati chain can be written as a system of first-order differential equations:
+$L^{m}u(z)+\sum_{j=1}^{m}\alpha_{j}(z)L^{j-1}u(z)+\alpha_{0}(z)=0,$
+where $m\in\mathbb{N}$, and the coefficient functions $\alpha_{j}(z)$ depend on $z$.
+Next, we introduce the system of differential equations corresponding to the Riccati chain.
+By defining $x_{k}\rightarrow L^{k-1}u$, the Riccati chain can be written as a system of first-order differential equations:
 $x_{k}^{\prime}$ $=x_{k+1}-cx_{1}x_{k},\quad$ $1\leq k\leq m-1,$ (2.1)
@@
 $x_{1}^{\prime}=-\alpha_{1}(z)x_{1}-\alpha_{0}(z)+cx_{1}^{2}.$
-The structure of the Riccati chain can be better understood by considering the weights of its terms.
-In particular, the weights of the terms $\alpha_{j}(z)$ are related to the scaling properties of the system.
-For example, the terms $\alpha_{j}(z)$ associated with the Riccati chain can be written as powers of $z$ with the following weights:
-[…]
-These weights play an essential role in understanding the quasi-homogeneous structure of the Riccati chain.
-Weight Classification for the Riccati Chain: The weights of the terms $\alpha_{j}(z)$ are crucial for classifying the Riccati chain’s solutions.
-Specifically, for the $m$-th Riccati chain given by (3.2), the weights are given by:
-$p_{k}=k\quad(1\leq k\leq m),\quad r\in\{1,\cdots,m+1\},\quad s=r+1.$
-Assuming the full vector field $F=(f_{1},\dots,f_{m})$ is quasi-homogeneous of degree +1+1, the coefficient functions $\alpha_{j}(z)$ satisfy the following weight classification:
+We assign weights to the variables $x_{k}$ and $z$ as follows:
+$\mathrm{wt}(x_{k})=p_{k}:=k,\quad 1\leq k\leq m,\quad\mathrm{wt}(z)=r\in\{1,\cdots,m+1\}$
+so that the vector field $f=(f_{1},\dots,f_{m})$ is quasi-homogeneous of degree +1+1, meaning that it satisfies the scaling relation:
+$f_{i}(\lambda^{p_{1}}x_{1},\dots,\lambda^{p_{m}}x_{m},\lambda^{r}z)=\lambda^{p_{i}+1}f_{i}(x_{1},\dots,x_{m},z),\quad i=1,\dots,m,\quad\lambda\in\mathbb{C}^{*}.$
+The coefficient functions $\alpha_{j}(z)$ satisfy the following weight classification:
 $\text{wt}(\alpha_{j})=m+1-j,\quad 0\leq j\leq m.$
@@
 $\alpha_{j}(z)=\begin{cases}A_{j}z^{\frac{m-j+1}{r}}&\text{if}\quad\frac{m-j+1}{r}\in\mathbb{Z},\\ 0&\text{otherwise}.\end{cases}$
-This weight classification helps in understanding the structure of the Riccati chain and its integrability.
-Examples for $n=1$
-For $n=1$, the Riccati chain reduces to the standard Riccati equation:
+Examples for $m=1$
+For $m=1$, the Riccati chain reduces to the standard Riccati equation:
 $x_{1}^{\prime}=-\alpha_{1}(z)x_{1}-\alpha_{0}(z)+cx_{1}^{2}.$
-The weight classification of this equation gives two types of quasi-homogeneous solutions:
+The weight classification of this equation gives two types of quasi-homogeneous equations:
 $x_{1}^{\prime}=-z+cx_{1}^{2}\quad\text{(Airy type)},$
@@
 Higher-order Examples and Chazy–Bureau Equations
-For higher-order cases, the Riccati chain leads to more complex equations. For $n=3$, the Riccati chain in scalar form reads:
+For higher-order cases, the Riccati chain leads to more complex equations. For $m=3$, the Riccati chain in scalar form reads:
 $u^{\prime\prime\prime}$ $+(\alpha_{3}(z)+4cu)u^{\prime\prime}+3c(u^{\prime})^{2}$
@@
 $u^{\prime\prime\prime}+3uu^{\prime\prime}+3(u^{\prime})^{2}+3u^{2}u^{\prime}=0.$
-For $n=4$, the Riccati chain yields the Fuchs-type equation:
+For $m=4$, the Riccati chain yields the Fuchs-type equation:
 $u^{(4)}+5uu^{\prime\prime\prime}+10u^{\prime}u^{\prime\prime}+10u^{2}u^{\prime\prime}+15u(u^{\prime})^{2}+10u^{3}u^{\prime}+u^{5}=0,$
@@
 3 Indicial Loci and Kovalevskaya Exponents
-We begin by recalling the autonomous part of the $m$-th Riccati chain(eq 3.2), written as an $m$-dimensional first-order system:
+We begin by considering the autonomous part of the $m$-th Riccati chain (2.1), written as an $m$-dimensional first-order system:
 $\frac{dx_{i}}{dz}=f_{i}^{A}(x),\quad i=1,\dots,m,$
@@
 $f_{m}^{A}(x)$ $=-cx_{1}x_{m},$
-and $c\in\mathbb{C}^{*}$ is a constant. We assign weights to the variables $x_{k}$ as follows:
-$\mathrm{wt}(x_{k})=p_{k}:=k,\quad 1\leq k\leq m,$
-so that the vector field $f^{A}=(f_{1}^{A},\dots,f_{m}^{A})$ is quasi-homogeneous of degree +1+1, meaning that it satisfies the scaling relation:
-$f_{i}^{A}(\lambda^{p_{1}}x_{1},\dots,\lambda^{p_{m}}x_{m})=\lambda^{p_{i}+1}f_{i}^{A}(x_{1},\dots,x_{m}),\quad i=1,\dots,m,\quad\lambda\in\mathbb{C}^{*}.$
-Next, we examine the formal Laurent series solutions of the truncated system of equations for small $z-z_{0}$. The solutions are of the form
+and $c\in\mathbb{C}^{*}$ is a constant.
+Next, we examine the formal Laurent series solutions of the truncated system of equations for small $z-z_{0}$. By quasihomogenity of the system, the solutions are of the form
 $x_{i}(z)=\xi_{i}(z-z_{0})^{-p_{i}}+\cdots,$
-where $p_{i}=i$ for $i=1,\dots,m$, and the leading coefficients $\xi=(\xi_{1},\dots,\xi_{m})$ satisfy the indicial equations
+where $p_{i}=i$ for $i=1,\dots,m$, and the leading coefficients $\xi=(\xi_{1},\dots,\xi_{m})$ are called indicial locus and satisfy the indicial equations
 $p_{i}\xi_{i}=f_{i}^{A}(\xi),\quad i=1,\dots,m.$ (3.2)
+Indicial loci
 Lemma 3.1.
@@
 Proof.
-Substituting $x_{i}(z)=\xi_{i}(z-z_{0})^{-p_{i}}$ into (4.1) and comparing the leading terms, we obtain the following recursive relations:
+Substituting $x_{i}(z)=\xi_{i}(z-z_{0})^{-i}$ into (3.2) and comparing the leading terms, we obtain the following recursive relations:
 $k\xi_{k}=\xi_{k+1}-c\xi_{1}\xi_{k},\quad 1\leq k\leq m-1,\quad-m\xi_{m}=-c\xi_{1}\xi_{m}.$
@@
 Thus, the only nonzero solutions to the indicial equations are the vectors $\xi^{(1)},\dots,\xi^{(m)}$, which correspond to the falling factorials $(n)_{k}/c$ for each $n$. ∎
-Kovalevskaya matrix and principal exponents
+Kovalevskaya matrix
 For a given indicial locus $\xi\in{\mathbb{C}}^{m}\setminus\{0\}$, the Kovalevskaya matrix is defined as
@@
 $\det(C_{m-n})=\prod_{k=n+1}^{m}(k-n-\lambda).$
-The matrix $A_{n}$ is lower triangular, and its entries are:
+The entries of matrix $A_{n}$ are:
 $A_{n}=\begin{pmatrix}1-2n-\lambda&1&0&\cdots&0\\ -(n)_{2}&2-n-\lambda&1&\cdots&0\\ -(n)_{3}&0&3-n-\lambda&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&1\\ -(n)_{n}&0&\cdots&0&-\lambda\end{pmatrix}$.
@@
 Thus, for each $n\in\{1,\dots,m\}$, the eigenvalues $\kappa=(\kappa_{1},\dots,\kappa_{m})$ of $K(\xi^{(n)})$ follow the pattern $(-n,-n+1,\dots,-1,1,\dots,m-n)$, which completes the proof. ∎
+Principal and lower indicial loci
+Let $\kappa(\xi^{k})$ denote the Kovalevskaya exponents associated with the indicial locus $\xi^{k}$ for $k\in\{1,\dots,m\}$. We can write the Kovalevskaya exponents of Riccati chain as following:
+Principal $\kappa(\xi^{1})=\{-1,1,\dots,m-1\},$
+Lower $\kappa(\xi^{2})=\{-2,-1,\dots,m-2\},$
+$\vdots$
+Lower $\kappa(\xi^{m-1})=\{-m+1,\dots,-1,1\},$
+Lower $\kappa(\xi^{m})=\{-m,\dots,-1\}.$
 Definition 3.3 (Principal and lower indicial loci).
-We call the indicial locus $\xi^{(1)}$ principal, since its K-exponents are exactly $\{-1,1,\dots,m-1\}$, i.e.
+We call the indicial locus $\xi$ is principal if the associated Laurent series solution exists and includes $m$ free parameters.
+If the number of nonnegative integer K-exponents are smaller than m1m-1, the locus is called an lower indicial locus.
+In the context of Kovalevskaya exponents of Riccati chain, the indicial locus $\xi^{(1)}$ is principal, since its K-exponents are exactly $\{-1,1,\dots,m-1\}$, i.e.
 one distinguished eigenvalue 1-1 and m1m-1 strictly positive integers.
-The loci $\xi^{(n)}$ with $m\geq 2$ are called lower indicial loci, in analogy with the terminology of the Painlevé test: they do not carry the full set of nonnegative integer Kovalevskaya exponents.
-5 Extended Painlevé Property
-The goal of this section is to prove that the Riccati chain satisfies the Extended Painlevé Property (EPP), as defined in [5]. For details on weighted projective spaces, readers can refer to [5].
-The Riccati hierarchy can be understood in the context of weighted projective spaces, which are spaces where the coordinates are scaled according to specific weights under a $\mathbb{C}^{*}$-action.
-More specifically, the weighted projective space $\mathbb{C}P^{m+2}(1,\cdots,m,m+1,m+2)$ associated with the $m$-th Riccati chain is defined by the equivalence relation on $\mathbb{C}^{4}\setminus\{0\}$:
-$(x_{1},\dots,x_{m},z,\varepsilon)\sim(\lambda x_{1},\dots,\lambda^{m}x_{m},\lambda^{m}z,\lambda^{m+1}\varepsilon),$
-where $\lambda\in\mathbb{C}^{*}$. This defines the weighted projective space, where each coordinate is rescaled according to a specific weight.
-The orbifold structure of this space is given by:
-$\mathbb{C}P^{m+1}(1,\cdots,m,r,s)=\mathbb{C}^{m+1}/\mathbb{Z}_{r}\cup\cdots\cup\mathbb{C}^{m+1}/\mathbb{Z}_{s},$
-where each $\mathbb{C}^{m+1}/\mathbb{Z}_{i}$ is a local chart for the projective space. The inhomogeneous coordinates on each chart are denoted by $(x_{1},\dots,x_{m},z)$.
-The transformation between the inhomogeneous coordinates $(x_{1},\dots,x_{m},z)$ on $\mathbb{C}^{m+1}/\mathbb{Z}_{s}$ and the coordinates $(X_{1},\dots,X_{j-1},X_{j+1},\dots,X_{m},Z,\varepsilon)$ on $\mathbb{C}^{m+1}/\mathbb{Z}_{j}$ is given by:
-$x_{i}=X_{i}\varepsilon^{-1/s}\quad(i\neq j),\quad x_{j}=\varepsilon^{-j/s},\quad z=Z\varepsilon^{-r/s},$
-where $\varepsilon$ is a local parameter for the action of $\mathbb{C}^{*}$.
-The Riccati Chain System in Local Coordinates
-The Riccati chain system on the local chart $\mathbb{C}^{m+1}/\mathbb{Z}_{j}$ can be written as the following system of differential equations:
-$\left\{\begin{array}[]{l}\displaystyle\frac{dX_{i}}{d\varepsilon}=\frac{1}{s\varepsilon}\left(iX_{i}-j\frac{f_{i}+\varepsilon G_{i}}{f_{j}+\varepsilon G_{j}}\right),\quad i=1,\dots,m,\quad i\neq j,\\ \displaystyle\frac{dZ}{d\varepsilon}=\frac{1}{s\varepsilon}\left(rZ-\frac{p_{j}\varepsilon}{f_{j}+\varepsilon G_{j}}\right),\end{array}\right$.
-(5.1)
-where $f_{i}=f_{i}(X_{1},\dots,X_{m},Z)$ are the components of the Riccati chain in the local coordinates.
-This system is rational and invariant under the $\mathbb{Z}_{j}$-action, reflecting the weighted projective space structure.
-For convenience, we rewrite the system (5.1) as an autonomous vector field in the form:
-$\left\{\begin{array}[]{l}\displaystyle\frac{dX_{i}}{dt}=iX_{i}-j\frac{X_{i+1}-cX_{1}X_{k}}{X_{j+1}-cX_{1}},\quad(i=1,\dots,m;i\neq j),\\ \displaystyle\frac{dZ}{dt}=rZ-\frac{j\varepsilon}{X_{j+1}-cX_{1}},\\ \displaystyle\frac{d\varepsilon}{dt}=s\varepsilon,\end{array}\right$.
-(5.2)
-where the new independent variable $t$ parametrizes the integral curves of the system.
-This formulation makes it easier to analyze the behavior of the Riccati chain in the context of weighted projective spaces.
-The following lemma describes the fixed points of the vector field in (5.2).
-Lemma 5.1.
-The point
-$\left(X_{1},\dots,X_{m},Z,\varepsilon\right)=\begin{cases}\left(\frac{(n)_{1}}{c}\left(\frac{(n)_{j}}{c}\right)^{-1/j},\dots,\frac{(n)_{j-1}}{c}\left(\frac{(n)_{j}}{c}\right)^{-(j-1)},0,\dots,0,0\right),&\text{for }j\leq n,\\ (0,\dots,0),&\text{for }j>n,\end{cases}$ (5.3)
-is a fixed point of the vector field in (5.2).
-Lemma 5.2.
-The eigenvalues of the Jacobi matrix of the vector field (5.2) at the fixed point (5.3) are given by $r$, $s$, and the Kovalevskaya exponents of the system (4.1), excluding the trivial exponent 1-1.
-This result has been established in [5].
-We now state the main result of this section:
+The loci $\xi^{(n)}$ with $n\geq 2$ are called lower indicial loci, since they do not carry the full set of nonnegative integer Kovalevskaya exponents.
+We further note that in the context of Riccati chains, negative exponents are determined by the leading coefficient on each branch.
+Remark 3.4.
+On each Riccati-chain Laurent branch determined by $\xi=\xi^{(n)}$, the branch parameter $\xi_{1}$ determines a negative Kovalevskaya exponent n-n via $n=-c\xi_{1}$.
+Corollary 3.5.
+For the Riccati chain at the indicial locus $\xi^{(n)}$, the vector $\xi^{(n)}$ itself is an eigenvector of $K(\xi^{(n)})$ with eigenvalue n-n, i.e.
+$K(\xi^{(n)})\,\xi^{(n)}=-n\,\xi^{(n)}.$
+Proof.
+In the explicit matrix form of $K(\xi^{(n)})-\lambda I$, set $\lambda=-n$.
+Since $(n)_{k}=0$ for $k>n$, the lower-right block does not interact with the first $n$ components.
+A direct substitution (using the same falling-factorial recursion as in the indicial equations) shows $(K(\xi^{(n)})+nI)\,\xi^{(n)}=0$.
+∎
+4 Commuting Vector Fields
+Let $F=(f_{1},\dots,f_{m})$ and $G=(g_{1},\dots,g_{m})$ be quasi-homogeneous polynomial vector fields on $\mathbb{C}^{m}$. We consider the following partial differential equations:
+$\frac{\partial x_{i}}{\partial z_{1}}=f_{i}(x),\quad\frac{\partial x_{i}}{\partial z_{2}}=g_{i}(x),\quad i=1,\dots,m,$
+where $x=(x_{1},\dots,x_{m})\in\mathbb{C}^{m}$ and $z_{1},z_{2}\in\mathbb{C}$. We assume the following conditions:
+(A1) $F$ and $G$ are quasi-homogeneous: there exist a tuple of positive integers $(a_{1},\dots,a_{m})\in\mathbb{N}^{m}$ and $\gamma\in\mathbb{N}$ such that
+$f_{i}(\lambda^{a_{1}}x_{1},\dots,\lambda^{a_{m}}x_{m})$ $=\lambda^{a_{i}+1}f_{i}(x_{1},\dots,x_{m}),$
+$g_{i}(\lambda^{a_{1}}x_{1},\dots,\lambda^{a_{m}}x_{m})$ $=\lambda^{a_{i}+\gamma}g_{i}(x_{1},\dots,x_{m}),$
+for any $\lambda\in\mathbb{C}$. We call $\gamma$ the degree of $G$ with respect to the weight $(a_{1},\dots,a_{m})$. The degree of $F$ is assumed to be 1.
+(A2) $F$ and $G$ commute with each other with respect to the Lie bracket:
+$F,G]=0,\quad\text{where}\quad[F,G]_{i}=\sum_{k=1}^{m}\left(f_{k}(x)\frac{\partial g_{i}}{\partial x_{k}}(x)-g_{k}(x)\frac{\partial f_{i}}{\partial x_{k}}(x)\right),\quad i=1,\dots,m.$
+(A3) $F(x)=0$ only when $x=0$.
+In the context of this paper, we consider the following proposition proposed in [16] related to the commutator and the indicial locus:
+Proposition 4.1.
+For a given indicial locus $c$, the identity $(K(c)+\gamma)G(c)=0$ holds. In particular, if $\gamma$ is not a K-exponent, then $G(c)=0$.
+We now consider the autonomous Riccati chain $F$ on $\mathbb{C}^{m}$, where each component $f_{k}$ is given by:
+$f_{k}(x)=x_{k+1}-cx_{1}x_{k},\quad 1\leq k\leq m-1,\quad f_{m}(x)=-cx_{1}x_{m},$
+with $c\in\mathbb{C}^{*}$ and $x=(x_{1},\dots,x_{m})\in\mathbb{C}^{m}$. The weights of the variables are assumed to be $\mathrm{wt}(x_{k})=k$ for $k=1,\dots,m$.
+Unique commuting vector field
 Theorem 4.2.
-The Riccati chain satisfies the Extended Painlevé Property (EPP) in the sense that it admits an mk+1m-k+1-parameter family of convergent Laurent series solutions, where the set $\{a_{i,n}\}$ includes mkm-k free parameters.
+Let $F$ be the autonomous Riccati chain with weights $\mathrm{wt}(x_{k})=k$.
+The only nontrivial quasi-homogeneous polynomial vector field commuting with $F$ is the unique commuting field $G$ of degree $\gamma=m$, given by
+$g_{i}=x_{i}x_{m}\quad\text{for}\quad i<m,\quad g_{m}=x_{m}^{2}.$
+For intermediate degrees $\gamma\in\{2,\dots,m-1\}$, no nontrivial commuting field exists.
 Proof.
-For each indicial locus $(\xi^{(n)})$, it corresponds to n1n-1 negative eigenvalues, mnm-n positive eigenvalues, and the trivial exponent 1-1.
-By Lemma 4.2, the eigenvalues of the Jacobi matrix at the fixed point are given by $\kappa$, and the parameters $r$ and $s$.
-By translation, the system can be decomposed into an mn+2m-n+2-dimensional unstable manifold and an n1n-1-dimensional stable manifold.
-We perform a suitable linear transformation, which allows us to rewrite the system in the form:
-$\left\{\begin{array}[]{l}\displaystyle\frac{d\mbox{$X$}_{u}}{dt}=J_{u}\mbox{$X$}_{u}+F_{u}(\mbox{$X$}_{u},\mbox{$X$}_{s}),\quad\mbox{$X$}_{u}\in\mathbb{C}^{m-n+2},\\ \displaystyle\frac{d\mbox{$X$}_{s}}{dt}=J_{s}\mbox{$X$}_{s}+F_{s}(\mbox{$X$}_{u},\mbox{$X$}_{s}),\quad\mbox{$X$}_{s}\in\mathbb{C}^{n-1}.\end{array}\right$.
-where $J_{u}$ and $J_{s}$ are the Jacobian matrices corresponding to the unstable and stable manifolds, respectively, with the real parts of the eigenvalues of $J_{u}$ being positive and those of $J_{s}$ being negative.
-By the unstable manifold theorem, there exists a local analytic function $\varphi$ such that $\varphi(0)=D\varphi(0)=0$, and the set $(\mbox{$X$}_{u},\varphi(\mbox{$X$}_{u}))$ forms the unstable manifold.
-The dynamics on the unstable manifold can then be expressed as:
-$\frac{d\mbox{$X$}_{u}}{dt}=J_{u}\mbox{$X$}_{u}+F_{u}(\mbox{$X$}_{u},\varphi(\mbox{$X$}_{u})).$
-Note that $J_{u}$ is semisimple and has distinct positive integer eigenvalues.
-Since for $j>n$, the fixed point is located at the origin, $(X_{1},\dots,X_{m},Z,\varepsilon)=(0,\dots,0,0)$, the fixed point lies at the origin.
-Furthermore, since the system does not contain resonant terms in the Poincaré domain, we can apply Poincaré’s linearization theory to linearize the unstable manifold, completing the proof that the Riccati chain satisfies the Extended Painlevé Property.
-∎
+We start by considering the quasi-homogeneous vector field $G=(g_{1},g_{2},\dots,g_{m})$, where each component $g_{i}$ is a polynomial of degree $i+\gamma$, which can be written in the general form:
+$g_{i}(x)=\sum_{j_{1}+2j_{2}+\dots+mj_{m}=i+\gamma}c_{j_{1},j_{2},\dots,j_{m}}x_{1}^{j_{1}}x_{2}^{j_{2}}\cdots x_{m}^{j_{m}},$
+where $c_{j_{1},j_{2},\dots,j_{m}}$ are constants, and the sum is over all multi-indices $(j_{1},j_{2},\dots,j_{m})$ such that the degree constraint $j_{1}+2j_{2}+\dots+mj_{m}=i+\gamma$ is satisfied.
+The commutator $F,G]_{m}(x)$ is given by:
+$F,G]_{m}(x)=\sum_{k=1}^{m}\left(f_{k}(x)\frac{\partial g_{m}}{\partial x_{k}}(x)-g_{k}(x)\frac{\partial f_{m}}{\partial x_{k}}(x)\right).$
+Substituting the expressions for $f_{k}(x)$ and $g_{k}(x)$, we have:
+$F,G]_{m}(x)=\sum_{k=1}^{m}\left(f_{k}(x)\frac{\partial g_{m}}{\partial x_{k}}(x)-g_{1}(x)(-cx_{m})-g_{m}(x)(-cx_{1})\right).$
+This simplifies to:
+$F,G]_{m}(x)=\sum_{k=1}^{m-1}(x_{k}-cx_{1}x_{k-1})\frac{\partial g_{m}}{\partial x_{k}}(x)-cx_{1}x_{m}\frac{\partial g_{m}}{\partial x_{m}}-g_{1}(x)(-cx_{m})-g_{m}(x)(-cx_{1}).$
+Now, we examine the critical term $(x_{k}-cx_{1}x_{k-1})\frac{\partial g_{m}}{\partial x_{k}}$.
+This term contains the factor $(x_{k}-cx_{1}x_{k-1})$, which introduces additional terms that cannot be canceled by any other terms in the expression.
+Specifically, when $k=m$, the structure of $g_{m}(x)=x_{m}^{2}$ creates terms that lead to an imbalance.
+Thus, we find that the commutator $F,G]_{m}(x)$ cannot vanish unless $g_{i}(x)=x_{i}x_{m}$ for $i=1,\dots,m-1$ and $g_{m}(x)=x_{m}^{2}$, which is the only solution that satisfies the commutation relation.
+Therefore, the only quasi-homogeneous polynomial vector fields commuting with $F$ are $F$ itself (degree 1) and the unique commuting field $G$ of degree $\gamma=m$, given by:
+$g_{i}(x)=x_{i}x_{m}\quad\text{for}\quad i=1,\dots,m-1,\quad g_{m}(x)=x_{m}^{2}.$
+For intermediate degrees $\gamma\in\{2,\dots,m-1\}$, no nontrivial commuting field exists. ∎
+Symmetry reduction
+The commuting vector field $G$ simplifies the Riccati chain by generating invariants along its flow.
+Specifically, along the flow of $G$, the quantities $y_{i}=\frac{x_{i}}{x_{m}}$ remain constant, which means the ratios of the $x_{i}$’s relative to $x_{m}$ are invariant.
+These invariants, interpreted as symmetries of the system, allow for a reduction in the dimensionality of the system.
+By exploiting these invariants, we reduce the Riccati chain to a simpler triangular system, where the equations are decoupled and can be solved sequentially.
+The general solution is then obtained in terms of rational functions of an arbitrary polynomial $u(z)$, where the solutions for $x_{1},x_{2},\dots,x_{m}$ are expressed as functions of $u(z)$.
+For the rest of section, we assume $c=1$ without loss of generality.
+Corollary 4.3.
+Let $F$ be the autonomous Riccati chain (3.1) with $c=1$, i.e.
+$x_{i}^{\prime}$ $=x_{i+1}-x_{1}x_{i},\qquad$ $1\leq i\leq m-1,$ (4.1)
+$x_{m}^{\prime}$ $=-x_{1}x_{m}.$
+Let $G$ be the commuting vector field of degree $\gamma=m$ given in Theorem 4.2,
+$g_{i}=x_{i}x_{m}\ (i<m),\qquad g_{m}=x_{m}^{2}.$
+On the open set $\{x_{m}\neq 0\}$, define
+$y_{i}:=\frac{x_{i}}{x_{m}}\quad(i=1,\dots,m-1),\qquad u:=\frac{1}{x_{m}}.$ (4.2)
+Then the Riccati chain (4.1) is transformed into the triangular system
+$y_{i}^{\prime}=y_{i+1}\quad(1\leq i\leq m-2),\qquad y_{m-1}^{\prime}=1,\qquad u^{\prime}=y_{1}.$ (4.3)
+Consequently, writing $t=z-z_{0}$ and introducing constants $a_{0},\dots,a_{m-1}\in{\mathbb{C}}$, one has
+$u(t)=\frac{t^{m}}{m!}+a_{m-1}\frac{t^{m-1}}{(m-1)!}+\cdots+a_{1}t+a_{0},$ (4.4)
+and
+$y_{i}(t)=u^{(i)}(t)\qquad(i=1,\dots,m-1).$ (4.5)
+The general solution of (4.1) is therefore
+$x_{m}(z)=\frac{1}{u(z-z_{0})},\qquad x_{i}(z)=\frac{y_{i}(z-z_{0})}{u(z-z_{0})}=\frac{u^{(i)}(z-z_{0})}{u(z-z_{0})}\quad(i=1,\dots,m-1).$ (4.6)
+Proof.
+Define $y_{i},u$ by (4.2). For $1\leq i\leq m-1$,
+$y_{i}^{\prime}=\left(\frac{x_{i}}{x_{m}}\right)^{\prime}=\frac{x_{i}^{\prime}x_{m}-x_{i}x_{m}^{\prime}}{x_{m}^{2}}.$
+Using (4.1), for $1\leq i\leq m-2$ we obtain
+$y_{i}^{\prime}=\frac{(x_{i+1}-x_{1}x_{i})x_{m}-x_{i}(-x_{1}x_{m})}{x_{m}^{2}}=\frac{x_{i+1}}{x_{m}}=y_{i+1},$
+and for i=m1i=m-1,
+$y_{m-1}^{\prime}=\frac{x_{m}}{x_{m}}=1.$
+Next,
+$u^{\prime}=\left(\frac{1}{x_{m}}\right)^{\prime}=-\frac{x_{m}^{\prime}}{x_{m}^{2}}=-\frac{-x_{1}x_{m}}{x_{m}^{2}}=\frac{x_{1}}{x_{m}}=y_{1}.$
+This proves (4.3).
+Integrating (4.3) yields $y_{m-1}(t)=t+a_{m-1}$, then $y_{m-2}(t)=\frac{t^{2}}{2!}+a_{m-1}t+a_{m-2}$, and inductively $y_{1}$ is a polynomial of degree m1m-1 whose coefficients are affine functions of the constants $a_{0},\dots,a_{m-1}$.
+$y_{m-k}(t)=\frac{t^{k}}{k!}+\frac{a_{m-1}t^{k-1}}{(k-1)!}+\frac{a_{m-2}t^{k-2}}{(k-2)!}+\cdots+a_{m-k},\quad k=1,\dots,m-1.$
+Finally, $u^{\prime}=y_{1}$ implies that $u$ is a polynomial of degree $m$, given by
+$u(t)=\frac{t^{m}}{m!}+\frac{a_{m-1}t^{m-1}}{(m-1)!}+\cdots+\frac{a_{1}t}{1!}+a_{0}.$
+The identities $y_{i}=u^{(i)}$ follow by differentiating $u^{\prime}=y_{1}$ repeatedly, hence (4.6) follows from $x_{m}=1/u$ and $x_{i}=y_{i}/u$. ∎
+In particulae, the general solution $x_{1}(z)$ is written as
+$x_{1}(z)$ $=\frac{y_{1}(z-z_{0})}{u(z-z_{0})}$ (4.7)
+$=\frac{u^{(1)}(z-z_{0})}{u(z-z_{0})}$
+$=\frac{\frac{t^{m-1}}{m-1!}+\frac{a_{m-1}t^{m-2}}{(m-2)!}+\cdots+a_{1}}{\frac{t^{m}}{m!}+\frac{a_{m-1}t^{m-1}}{(m-1)!}+\cdots+\frac{a_{1}t}{1!}+a_{0}}$
+$=\frac{mt^{m-1}+m(m-1)a_{m-1}t^{m-2}+\cdots+m!a_{1}}{t^{m}+ma_{m-1}t^{m-1}+\cdots+m!a_{0}}$
+$=\frac{mt^{m-1}+(m-1)b_{m-1}t^{m-2}+\cdots+(m-1)!b_{1}}{t^{m}+b_{m-1}t^{m-1}+\cdots+b_{0}}$
+where $b_{m-k}=m(m-1)\cdots(m-k+1)a_{m-k}$.
+Blow-up method
+In the context of the 2-dimensional Riccati chain, we apply the blow-up method commonly used in Painlevé analysis to resolve singularities[13, 14].
+This procedure provides a regular foliation on a smooth model and will be shown to yield the same general solution as derived through symmetry reduction.
+The blow-up method introduces a sequence of transformations that compactify the space and resolve indeterminacies at infinity, leading to a simpler structure that can be analyzed more easily.
+Compactification and a base point at infinity
+Consider the $2$-dimensional autonomous system
+$x_{1}^{\prime}=x_{2}-x_{1}^{2},\qquad x_{2}^{\prime}=-x_{1}x_{2}.$ (4.8)
+On $\mathbb{P}$$\times{\mathbb{P}}$ introduce the chart at infinity
+$(X,Y):=\Bigl(\frac{1}{x_{1}},\frac{1}{x_{2}}\Bigr),$
+so that the induced vector field is
+$X^{\prime}=1-\frac{X^{2}}{Y},\qquad Y^{\prime}=\frac{Y}{X}.$
+The point $(X,Y)=(0,0)$ is a base point, since the right-hand sides are indeterminate there.
+Blow-up
+Blow up $(X,Y)=(0,0)$. In the chart X=u,Y=uvX=u,\ Y=uv one obtains
+$u^{\prime}=\frac{v-u}{v},\qquad v^{\prime}=1.$ (4.9)
+Equation (4.9) is integrable:
+$v(z)=z-z_{0},\qquad u(z)=\frac{(z-z_{0})^{2}+2C}{2(z-z_{0})}=\frac{z^{2}-2zz_{0}+z_{0}^{2}+2C}{2z-2z_{0}},$
+where $C$ is an arbitrary constant.
+With $b_{1}=2z_{0}$ and $b_{2}=z_{0}^{2}+2C$, we recover the same general solution as (4.7) after transforming back to $(x_{1},x_{2})$.
+5 Annular Laurent series expansions
+In this section, we examine the Laurent expansion of the solution $x_{1}(z)$ in the annular regions of the complex plane, and explore how the negative Kovalevskaya exponent n-n varies with the choice of region (disk, annulus, or exterior) for the expansion.
+We begin by rewritting our general solution (4.7) into a sum over simple poles:
+$x_{1}(z)=\sum_{j=1}^{m}\frac{1}{z-z_{j}},$ (5.1)
+where $z_{1},\dots,z_{m}$ are the singularities of the solution.
+Next, we introduce the transformation $w:=z-z_{1}$ and $\delta_{j}:=z_{j}-z_{1}$ for $j\geq 2$, and rewrite the solution as:
+$x_{1}(w)=\frac{1}{w}+\sum_{j=2}^{m}\frac{1}{w-\delta_{j}}.$ (5.2)
+For each term $\frac{1}{w-\delta_{j}}$, we examine the geometric series expansions depending on the region where the expansion is performed.
+Geometric series expansion
+We consider the following two cases based on the relative magnitudes of $|w|$ and $|\delta_{j}|$:
+  • Inside the disc ($|w|<|\delta_{j}|$): In this case, the geometric series expansion is:
+    $\frac{1}{w-\delta_{j}}=-\frac{1}{\delta_{j}}\cdot\frac{1}{1-w/\delta_{j}}=-\sum_{k=0}^{\infty}\frac{w^{k}}{\delta_{j}^{k+1}},$ (5.3)
+    which does not contain a $w^{-1}$ term.
+  • Outside the disc ($|w|>|\delta_{j}|$): Here, the expansion becomes:
+    $\frac{1}{w-\delta_{j}}=\frac{1}{w}\cdot\frac{1}{1-\delta_{j}/w}=\frac{1}{w}\sum_{k=0}^{\infty}\Big(\frac{\delta_{j}}{w}\Big)^{k},$ (5.4)
+    where the leading term is $1/w$.
+Annular Regions
+Let $0<r_{2}\leq r_{3}\leq\cdots\leq r_{m}$ be the ordered moduli of $\delta_{j}$ (i.e. the radii $|\delta_{j}|$ in nondecreasing order). Define annular regions
+$A_{n}:=\{w\in{\mathbb{C}}:\ r_{n}<|w|<r_{n+1}\},\qquad n=1,\dots,m,$ (5.5)
+with the conventions $r_{1}:=0$ and $r_{m+1}:=+\infty$. Thus: $A_{1}$ is a punctured disc around $w=0$, intermediate $A_{n}$ are annuli, and $A_{m}$ is the exterior region.
+For a fixed annulus $A_{n}$, split the set of poles other than $z_{1}$ into
+$I_{\rm in}(n):=\{\,j\in\{2,\dots,m\}:\ |\delta_{j}|<|w|\ \text{for }w\in A_{n}\,\},\qquad I_{\rm out}(n):=\{2,\dots,m\}\setminus I_{\rm in}(n).$
+By construction, $|I_{\rm in}(n)|=n-1$ and $|I_{\rm out}(n)|=m-n$.
+Proposition 5.1.
+Fix $n\in\{1,\dots,m\}$ and an annulus $A_{n}$ as in (5.5). Then $x_{1}(w)$ admits on $A_{n}$ a Laurent expansion of the form
+$x_{1}(w)=\frac{n}{w}+H_{n}(w),$ (5.6)
+where $H_{n}(w)$ is holomorphic on $A_{n}$ and admits a Laurent series that converges normally on $A_{n}$. Specifically, the explicit form of $H_{n}(w)$ is:
+$H_{n}(w)=\sum_{j\in I_{\rm in}(n)}\sum_{k=1}^{\infty}\frac{\delta_{j}^{k}}{w^{k+1}}\;-\;\sum_{j\in I_{\rm out}(n)}\sum_{k=0}^{\infty}\frac{w^{k}}{\delta_{j}^{k+1}}.$ (5.7)
+In particular, the coefficient of $w^{-1}$ in this Laurent expansion is $n$.
+Proof.
+We begin by recalling the expression for $x_{1}(w)$ given in (5.2).
+For each $j\in I_{\rm in}(n)$, we have $|\delta_{j}/w|<1$, so the series expansion (5.3) converges absolutely and uniformly on compact subsets of $A_{n}$.
+Similarly, for each $j\in I_{\rm out}(n)$, we have $|w/\delta_{j}|<1$, so the series (5.4) converges absolutely and uniformly on compact subsets of $A_{n}$.
+Substituting these expansions into (5.2) yields:
+$x_{1}(w)=\frac{1}{w}+\sum_{j\in I_{\rm in}(n)}\left(\frac{1}{w}+\sum_{k\geq 1}\frac{\delta_{j}^{k}}{w^{k+1}}\right)+\sum_{j\in I_{\rm out}(n)}\left(-\sum_{k\geq 0}\frac{w^{k}}{\delta_{j}^{k+1}}\right).$
+Thus, the coefficient of $w^{-1}$ is computed as:
+$1+|I_{\rm in}(n)|=1+(n-1)=n.$
+The remaining terms form the Laurent series $H_{n}(w)$, which is holomorphic on $A_{n}$ and has the explicit form:
+$H_{n}(w)=\sum_{j\in I_{\rm in}(n)}\sum_{k=1}^{\infty}\frac{\delta_{j}^{k}}{w^{k+1}}-\sum_{j\in I_{\rm out}(n)}\sum_{k=0}^{\infty}\frac{w^{k}}{\delta_{j}^{k+1}}.$
+Remark 5.2.
+By Proposition 5.1, the negative Kovalevskaya exponent can be interpreted as the annulus index (disk/annulus/exterior), which indicates the residue of $x_{1}$ at the movable pole.
+For $m\geq 3$, there are $m$ regions $A_{n}$, and thus multiple admissible Laurent expansions around a given movable pole.
+The inner expansion corresponds to $A_{1}$ (the punctured disk), the outer expansion corresponds to $A_{m}$ (the exterior region), and the intermediate annuli $A_{n}$ provide Laurent expansions with contributions from both inner and outer regions, with the leading coefficient being $n\in\{1,\dots,m\}$.
 Acknowledgments. The author wishes to express sincere gratitude to Professor Hayato Chiba for his valuable guidance and instruction.